# Revisiting Isocurvature Bounds on the Minimal QCD Axion

### v1

@@ 4.2 Numerical solution across the parameter space
We solve numerically Eqs. 4.4a and 4.7b in the background evolution of the inflaton during the observable epoch of inflation and the preheating phase. For our illustrative purposes, we consider two inflationary models (described in detail in Appendix B).
@@
  • We focus on the pre-inflationary parameter space, requiring then from Eq. 2.8 that $f_{a}>T_{\text{max}}$. For each inflationary model, the ratio between $T_{\text{max}}$ and the temperature corresponding to an instantaneous reheating $T_{\mathrm{inst.\,RH}}\equiv(90H_{I}^{2}M_{\textsc{p}}^{2}/(\pi^{2}g_{\star}))^{1/4}$ is fixed, so
[…]
@@
  • The term $2\xi R|\Phi|^{2}$ in Eq. 3.2 corresponds to a negative contribution to the total energy density during inflation, which should not overcome the inflaton one. This amounts to
@@
    $\left|V\left(\langle\Phi\rangle=\tfrac{1}{\sqrt{2}}f_{a}^{\text{(inf)}}\right)\right|\approx\frac{\xi^{2}(12H_{I}^{2})^{2}}{\lambda_{\Phi}}<3H_{I}^{2}M_{\textsc{p}}^{2}\,.$ (4.9)
@@
  • The coefficient $\xi$ can be bounded from above by unitarity arguments. In the scattering $\Phi\Phi\to\Phi\Phi$ mediated by a graviton exchange, the operator $2\xi R|\Phi|^{2}$ contributes with a series of couplings of $\Phi\Phi$ to gravitons. The leading term comes at mass dimension 5 [85, 86], with an amplitude of order $\xi^{2}E^{2}/M_{\textsc{p}}^{2}$, where $E$ is the energy of the process. We impose that this amplitude is smaller than 1 at the highest energy scale of the problem, the inflationary scale $(H_{I}M_{\textsc{p}})^{1/2}$.
@@
  • The smallness of the quartic coupling $\lambda_{\Phi}$ that we consider means we should check the size of quantum corrections (see also [87, 88]). In the model described by Eq. 3.2, the argument of dimensional analysis (by restoring $\hbar$ to make couplings dimensionful [89]) shows that the only terms which might correct the quartic coupling at loop order $\hbar/(4\pi)^{2}$ are $\propto\xi R$ or $\propto m_{\Phi}^{2}(=\lambda_{\Phi}f_{a}^{2})$. Any correction to $\lambda_{\Phi}$ would then be suppressed by the second power of one of them over $f_{a}^{4}$ or $M_{\textsc{p}}^{4}$. All of these combinations are small enough in our sample. Or of course the loop correction could be proportional to $\lambda_{\Phi}$ itself, but then is necessarily small enough. We comment in Section A.2 on an analogous requirement in case the saxion couples à la KSVZ via a Yukawa coupling to coloured fermions. To be thorough, we should also estimate gravitational loop corrections to $\lambda_{\Phi}$. Some of these graviton loops may be quartically divergent. The maximum contribution these can make is $\sim\frac{1}{(4\pi)^{2}}\frac{\text{cutoff}^{4}}{M_{\textsc{p}}^{4}}$. To be conservative, if we take the cutoff of the gravitational EFT to be as high as $f_{a}$ we still find that this contribution is negligible in most of parameter space and would require a mild fine-tuning of $\lambda_{\Phi}$ for the points in the lower-right corner of Fig. 10. Even this may be an overestimate since whatever UV physics cuts off these graviton loops, for example supersymmetry, could come in at a scale well below $f_{a}$. Thus it seems likely that gravitational corrections do not give too large a contribution to $\lambda_{\Phi}$ and do not require fine-tuning.
@@
We perform a scan of the 4-dimensional parameter space $(H_{I},f_{a},\xi,\lambda_{\Phi})$ as follows. We draw a random sample of points satisfying the isocurvature bound Eq. 2.6 with $f_{a}^{\text{(inf)}}$ defined as in Eq. 3.4, and further impose the constraints described in the previous bullet points. The conditions that we impose on the sample are collected in Eqs. (4.11a)-(4.11f).
@@
Then, we compute the evolution of the saxion background $\overline{X}(t)$ in this background, using Eq. 4.4a. The initial condition for $\overline{X}$ is not important, as it quickly relaxes to its minimum during inflation.
Finally, we solve the linearised equations of motion for the perturbations $y_{k}$ (Eq. 4.7b) for a range of axion modes with physical momentum $k\sim\mathcal{O}(0.1-10)k_{e}$, with $k_{e}=a_{e}H_{e}$ crossing the Hubble radius at the end of inflation. The initial condition for each mode $y_{k}$ is set by the quantum fluctuations of the Bunch-Davies vacuum, which at the time $t_{\crossproduct}$ when the mode crosses the Hubble radius reads $|y_{k}(t_{\crossproduct})|\approx H_{\crossproduct}/\sqrt{2k^{3}}$. We perform this numerical calculation for enough time to detect whether the parametric resonance induced on the axion is strong enough to amplify its power spectrum more than $(\pi f_{a})^{2}$ and restore $U(1)_{\textsc{pq}}$. We compute the power spectrum (at a late time when $\overline{X}(t)\approx f_{a}$ and $Y(x)\approx a(x)$) as $\langle(a(t,x)-a_{i})^{2}\rangle\approx\langle Y(t,x)^{2}\rangle=\int|y_{k}(t)|^{2}\mathrm{d}^{3}k$.
@@
The result of this scan of the parameter space is shown in Fig. 3 as projected on the $(H_{I},f_{a})$ plane. The colour code marks the enhancement of axion fluctuations as a result of the dynamics after the end of inflation. Points in blue experience little or no parametric resonance, whereas the evolution of light blue and rose points amplifies the fluctuations in the misalignment angle to the point where our linearised analysis breaks down and $U(1)_{\textsc{pq}}$ is non-thermally restored, leading to the overproduction of axions at late times (in this part of parameter space). More details about these results are given in Section A.1 and Figs. 8 and 9.
Figure 4: Growth of the axion fluctuations due to parametric resonance as a function of the enhancement of $f_{a}$ during inflation, $f_{a}^{\text{(inf)}}\approx\sqrt{24\xi/\lambda_{\Phi}}H_{I}$, for our scenario of warm inflation. The red horizontal line marks $f_{a}^{\text{(inf)}}/f_{a}\approx 700$, which roughly divides the points featuring inefficient or short parametric resonance (in blue) and string formation (light orange). Points which display an amplification $\tfrac{1}{\pi}\delta\theta_{\text{(par.res.)}}<10^{-9}$, or $>10^{2}$, are all marked on the respective threshold. Some points on the lower left align on diagonal lines because of our sampling of integer powers of $\xi/\lambda_{\Phi}$.
@@
$\frac{1}{\pi}\delta\theta_{\text{(par.res.)}}\equiv\frac{\langle a(x)^{2}\rangle^{1/2}}{\pi f_{a}}$ (4.10)
versus the enhancement of $f_{a}$ during inflation $f_{a}^{\text{(inf)}}/f_{a}\approx\sqrt{24\xi/\lambda_{\Phi}}H_{I}/f_{a}$. The result in Fig. 4 shows that the enhancement $f_{a}^{\text{(inf)}}/f_{a}$ is a very good proxy to predict the strength of the subsequent parametric resonance. The origin of this correlation is clear, as the linearised Eq. 4.7b for the perturbations $y_{k}$ has a non-linear driving source proportional to the displacement of $\overline{X}(t)$ from the minimum $f_{a}^{\text{(inf)}}(t)$. The larger is the excursion of $f_{a}^{\text{(inf)}}$ at the end of inflation when moving to $f_{a}$, the longer is the duration of the parametric resonance for $y_{k}$. Section A.1 discusses this point with further details.
As a result, we can posit an approximate condition to ensure that $U(1)_{\textsc{pq}}$ is not restored due to parametric resonance: for the model of warm inflation considered in this section, this condition is $f_{a}^{\text{(inf)}}/f_{a}\lesssim 700$. As clear in Fig. 4, when this condition is satisfied, the amplification $\delta\theta_{\text{(par.res.)}}$ of the power spectrum of the misalignment angle remains $<\mathcal{O}(10^{-3})$, ensuring the validity of our linearised analysis. This condition is model dependent: the details of the saxion-inflaton coupling and of the model for inflation and reheating, affect this quantitative threshold. Depending on the complexity of the dynamics, and on the possibility (for cold inflation) that the oscillations of the inflaton at preheating might act as additional source of parametric resonance, this condition to prevent string formation might be a bit less sharp than in Fig. 4. At any rate, also for the models of cold inflation that we consider in Appendix B we find that the condition $f_{a}^{\text{(inf)}}/f_{a}\lesssim\mathcal{O}(100)$ is an excellent proxy to avoid the formation of axion strings.
@@
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\left(\frac{90M_{\textsc{p}}^{2}}{\pi^{2}g_{\star}}\right)^{1/4}\varepsilon_{i}}$ $<\frac{f_{a}}{H_{I}^{1/2}}$ avoidance of PQ restoration at reheating, Eq. 4.8 (4.11a)
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\frac{2.0\cdot 10^{-6}}{\,\text{GeV}^{7/5}}}$ $>\frac{f_{a}}{H_{I}^{12/5}}$ standard isocurvature for axion DM, Eqs. 2.6 and 2.7 (4.11b)
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\frac{\,\text{GeV}^{7/6}}{7.5\cdot 10^{-7}}}\frac{\xi}{\lambda_{\Phi}}$ $>f_{a}^{7/6}$ isocurvature bound for axion DM with $f_{a}^{\text{(inf)}}$, Eq. 3.4 (4.11c)
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\frac{4\sqrt{3}}{M_{\textsc{p}}}}\frac{\xi}{\lambda_{\Phi}^{1/2}}$ $<\frac{1}{H_{I}}$ saxion energy density not exceeding $\rho_{\text{tot}}$, Eq. 4.9 (4.11d)
$\frac{1}{{\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}M_{\textsc{p}}^{1/2}}}\xi$ $<\frac{1}{H_{I}^{1/2}}$ unitarity bound on $\xi$ from graviton-mediated $\Phi\Phi\to\Phi\Phi$ (4.11e)
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\frac{4\sqrt{6}}{M_{\textsc{p}}}}\frac{\xi}{\lambda_{\Phi}^{1/2}}$ $\lesssim\frac{1}{H_{I}}$ EFT validity ($2\xi(f_{a}^{\text{(inf)}})^{2}R$ sub-leading wrt $\tfrac{1}{2}M_{\textsc{p}}^{2}R$) (4.11f)
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\mathcal{O}\left(10^{-2}\right)}\sqrt{\frac{\xi}{\lambda_{\Phi}}}$ $\lesssim\frac{f_{a}}{H_{I}}$ avoiding string formation (Fig. 4) (4.11g)
All of these conditions are set analytically except the last condition Eq. 4.11g which is an approximate result of our numerical analysis. The numerical coefficient of the last condition would be $\approx\sqrt{24}/700$ in the warm inflation model we consider but is different in other models, though in general is $\sim 10^{-2}$.
@@
Being equipped with Eq. 4.11g to determine the formation of axion strings, we can understand analytically the results of the scatter plot in Fig. 3. Let us fix the ratio $\xi/\lambda_{\Phi}$, or equivalently $f_{a}^{\text{(inf)}}/H_{I}$. For a fixed value there is an allowed wedge in the plane $(H_{I},f_{a})$, as marked in Fig. 5 for two choices of $f_{a}^{\text{(inf)}}/H_{I}$ (the blue and green regions). The isocurvature bound with $f_{a}^{\text{(inf)}}$, Eq. 4.11c, gives the horizontal upper limit of these regions. The condition Eq. 4.11g to prevent string formation translates into a diagonal upper limit on $H_{I}$ as seen in the right hand sides of the blue and green regions in Fig. 5. As we vary $f_{a}^{\text{(inf)}}/H_{I}$, the envelope of these wedges identifies the boundary marked in orange in Fig. 5. If we denote by $R_{f_{a}}$ the upper limit on $f_{a}^{\text{(inf)}}/f_{a}$ identified in Fig. 4 (in that case, $R_{f_{a}}=700$), this right-hand side boundary of the parameter space “rescued” from isocurvature reads (using Eq. 4.11c for the isocurvature bound)
[…]
Finally, there is an upper limit on $f_{a}$ that can be derived by the combination of Eqs. 4.11f and 4.11c. This bound reads
[…]
and is marked in purple in Fig. 5, with a choice of $\lambda_{\Phi}=10^{-16}$ for the figures of the paper (a benchmark choice which selects most of our blue points in Fig. 10). This line also roughly coincides with an anaologous combination of the isocurvature constraint Eq. 4.11c and the requirement that the saxion VEV $f_{a}^{\text{(inf)}}$ during inflation does not reach Planckian values.
@@
  • We neglect the fluctuations $x_{k}$ of the saxion, which display a similar growth as the axion fluctuations and are then irrelevant for us, as long as $\tfrac{1}{\pi}\delta\theta_{\text{(par.res.)}}$ is safely $\ll 1$. One should consider also that the saxion, while passing over the wine bottle potential across the origin, might stop in the angle $\theta_{i}$ (as in the top-left panel of 9) or in $\theta_{i}+\pi$ (as in Fig. 8). This only affects the initial value of the misalignment angle, which we assume to be the right one to yield the axion as DM. We do not need to worry instead that the final angular position ($\theta_{\text{i}}$ or $\theta_{\text{i}}+\pi$) might vary across different spatial patches, depending on the local quantum fluctuations of the saxion. The reason is that, for the “rescued” region of Fig. 6, the saxion crosses the origin less than 10 times, and in that region the relative quantum fluctuations of the saxion $(H_{I}/(2\pi))/f_{a}$ are much smaller than $1/10$, implying that the saxion stops in the same field position across the whole observable universe.
@@
    The decay of the saxion depends on the content of the PQ sector. A model-independent channel is the decay to axions, proportional to $m_{\Phi}^{3}/\langle\phi\rangle^{2}\sim\lambda_{\Phi}^{3/2}f_{a}$ [39]. Another channel that is generically present is the decay to SM gluons, also proportional to $m_{\Phi}^{3}/\langle\phi\rangle^{2}$. Other decay channels, as e.g. to KSVZ fermions (when kinematically allowed, see Section A.2), can deplete the saxion energy density into the SM plasma. For the “safe” blue points of the scan shown in Fig. 3, we find that the decays to axions or gluons happen around a similar temperature (due to the same parametric scaling of their decay rates) around $T\sim 10^{2-9}\,\text{GeV}$.
    For about half of the points, the decay occurs after the domination by the saxion of the energy budget of the Universe. In this case, we might worry about isocurvature perturbations on CMB scales. Saxion perturbations on large scales are tiny at early times (the saxion being a massive degree of freedom during inflation [90, 91, 92]). They inherit the adiabatic perturbations on large scales from the spatial dependence of the Hubble rate: there is effectively a single clock for curvature perturbations. We find then no general issue with the saxion inducing a matter-dominated epoch, given that it decays way before BBN in the parameter space of interest to us.

### v2

@@ 4.2 Numerical solution across the parameter space
We solve numerically Eqs. 16a and 19b in the background evolution of the inflaton during the observable epoch of inflation and the preheating phase. For our illustrative purposes, we consider two inflationary models (described in detail in Appendix B).
@@
In the main text we will give the results for the warm inflation model. This warm inflation model yields a slightly larger region of parameter space which is safe from isocurvature compared to the particular cold inflation model we used. This is because our chosen warm inflation model ends more smoothly, so the saxion relaxes towards its ultimate late-time minimum more gradually than in the cold inflation model we chose, where the end is more abrupt. The more abrupt end will tend to drive somewhat more parametric resonance production of the axion. The results for the particular cold inflation model we chose are presented in Appendix B (Fig. 13). We have not done a scan over all possible inflation models (nor over other possible UV realizations of the axion/saxion). We just chose two inflation models to illustrate the model-dependence. So in particular we certainly cannot state a theorem that warm inflation will in general give a larger parameter space than cold inflation.
@@
  • We focus on the pre-inflationary parameter space, requiring then from Eq. 8 that $f_{a}>T_{\text{max}}$. For each inflationary model, the ratio between $T_{\text{max}}$ and the temperature corresponding to an instantaneous reheating $T_{\mathrm{inst.\,RH}}\equiv(90H_{I}^{2}M_{\textsc{p}}^{2}/(\pi^{2}g_{\star}))^{1/4}$ is fixed, so
    $T_{\text{max}}=8.4\cdot 10^{13}\,\text{GeV}\left(\frac{H_{I}}{10^{10}\,\text{GeV}}\right)^{1/2}\cdot\varepsilon_{i}\,,\quad\varepsilon_{i}=0.03-0.32\text{ for our infl.\,models}.$ (20)
@@
  • The term $2\xi R|\Phi|^{2}$ in Eq. 10 corresponds to a negative contribution to the total energy density during inflation, which should not overcome the inflaton one. This amounts to
@@
    $\left|V\left(\langle\Phi\rangle=\tfrac{1}{\sqrt{2}}f_{a}^{\text{(inf)}}\right)\right|\approx\frac{\xi^{2}(12H_{I}^{2})^{2}}{\lambda_{\Phi}}<3H_{I}^{2}M_{\textsc{p}}^{2}\,.$ (21)
@@
  • The coefficient $\xi$ can be bounded from above by unitarity arguments. In the scattering $\Phi\Phi\to\Phi\Phi$ mediated by a graviton exchange, the operator $2\xi R|\Phi|^{2}$ contributes with a series of couplings of $\Phi\Phi$ to gravitons. The leading term comes at mass dimension 5 Burgess:2009ea; Barbon:2009ya, with an amplitude of order $\xi^{2}E^{2}/M_{\textsc{p}}^{2}$, where $E$ is the energy of the process. We impose that this amplitude is smaller than 1 at the highest energy scale of the problem, the inflationary scale $(H_{I}M_{\textsc{p}})^{1/2}$.
@@
  • The smallness of the quartic coupling $\lambda_{\Phi}$ that we consider means we should check the size of quantum corrections (see also Markkanen:2018bfx; Kozow:2022whq). In the model described by Eq. 10, the argument of dimensional analysis (by restoring $\hbar$ to make couplings dimensionful Panico:2015jxa) shows that the only terms which might correct the quartic coupling at loop order $\hbar/(4\pi)^{2}$ are $\propto\xi R$ or $\propto m_{\Phi}^{2}(=\lambda_{\Phi}f_{a}^{2})$. Any correction to $\lambda_{\Phi}$ would then be suppressed by the second power of one of them over $f_{a}^{4}$ or $M_{\textsc{p}}^{4}$. All of these combinations are small enough in our sample. Or of course the loop correction could be proportional to $\lambda_{\Phi}$ itself, but then is necessarily small enough. We comment in Section A.2 on an analogous requirement in case the saxion couples à la KSVZ via a Yukawa coupling to coloured fermions. To be thorough, we should also estimate gravitational loop corrections to $\lambda_{\Phi}$. Some of these graviton loops may be quartically divergent. The maximum contribution these can make is $\sim\frac{1}{(4\pi)^{2}}\frac{\text{cutoff}^{4}}{M_{\textsc{p}}^{4}}$. To be conservative, if we take the cutoff of the gravitational EFT to be as high as $f_{a}$ we still find that this contribution is negligible in most of parameter space and would require a mild fine-tuning of $\lambda_{\Phi}$ for the points in the lower-right corner of Fig. 7. Even this may be an overestimate since whatever UV physics cuts off these graviton loops, for example supersymmetry, could come in at a scale well below $f_{a}$. Thus it seems likely that gravitational corrections do not give too large a contribution to $\lambda_{\Phi}$ and do not require fine-tuning.
@@
We perform a scan of the 4-dimensional parameter space $(H_{I},f_{a},\xi,\lambda_{\Phi})$ as follows. We draw a random sample of points satisfying the isocurvature bound Eq. 6 with $f_{a}^{\text{(inf)}}$ defined as in Eq. 12, and further impose the constraints described in the previous bullet points. The conditions that we impose on the sample are collected in Eqs. (23a)-(23f).
@@
Then, we compute the evolution of the saxion background $\overline{X}(t)$ in this background, using Eq. 16a. The initial condition for $\overline{X}$ is not important, as it quickly relaxes to its minimum during inflation.
Finally, we solve the linearised equations of motion for the perturbations $y_{k}$ (Eq. 19b) for a range of axion modes with physical momentum $k\sim\mathcal{O}(0.1-10)k_{e}$, with $k_{e}=a_{e}H_{e}$ crossing the Hubble radius at the end of inflation. The initial condition for each mode $y_{k}$ is set by the quantum fluctuations of the Bunch-Davies vacuum, which at the time $t_{\crossproduct}$ when the mode crosses the Hubble radius reads $|y_{k}(t_{\crossproduct})|\approx H_{\crossproduct}/\sqrt{2k^{3}}$. We perform this numerical calculation for enough time to detect whether the parametric resonance induced on the axion is strong enough to amplify its power spectrum more than $(\pi f_{a})^{2}$ and restore $U(1)_{\textsc{pq}}$. We compute the power spectrum (at a late time when $\overline{X}(t)\approx f_{a}$ and $Y(x)\approx a(x)$) as $\langle(a(t,x)-a_{i})^{2}\rangle\approx\langle Y(t,x)^{2}\rangle=\int|y_{k}(t)|^{2}\mathrm{d}^{3}k$.
@@
The result of this scan of the parameter space is shown in Fig. 3 as projected on the $(H_{I},f_{a})$ plane. The colour code marks the enhancement of axion fluctuations as a result of the dynamics after the end of inflation. Points in blue experience little or no parametric resonance, whereas the evolution of light blue and rose points amplifies the fluctuations in the misalignment angle to the point where our linearised analysis breaks down and $U(1)_{\textsc{pq}}$ is non-thermally restored, leading to the overproduction of axions at late times (in this part of parameter space). More details about these results are given in Section A.1 and Figs. 9 and 10.
Figure 4: Growth of the axion fluctuations due to parametric resonance as a function of the enhancement of $f_{a}$ during inflation, $f_{a}^{\text{(inf)}}\approx\sqrt{24\xi/\lambda_{\Phi}}H_{I}$, for our scenario of warm inflation. The red horizontal line marks $f_{a}^{\text{(inf)}}/f_{a}\approx 700$, which roughly divides the points featuring inefficient or short parametric resonance (in blue) and string formation (light orange). Points which display an amplification $\tfrac{1}{\pi}\delta\theta_{\text{(par.res.)}}<10^{-9}$, or $10^{2}$, are all marked on the respective threshold. Some points on the lower left align on diagonal lines because of our sampling of integer powers of $\xi/\lambda_{\Phi}$.
@@
$\frac{1}{\pi}\delta\theta_{\text{(par.res.)}}\equiv\frac{\langle a(x)^{2}\rangle^{1/2}}{\pi f_{a}}$ (22)
versus the enhancement of $f_{a}$ during inflation $f_{a}^{\text{(inf)}}/f_{a}\approx\sqrt{24\xi/\lambda_{\Phi}}H_{I}/f_{a}$. The result in Fig. 4 shows that the enhancement $f_{a}^{\text{(inf)}}/f_{a}$ is a very good proxy to predict the strength of the subsequent parametric resonance. The origin of this correlation is clear, as the linearised Eq. 19b for the perturbations $y_{k}$ has a non-linear driving source proportional to the displacement of $\overline{X}(t)$ from the minimum $f_{a}^{\text{(inf)}}(t)$. The larger is the excursion of $f_{a}^{\text{(inf)}}$ at the end of inflation when moving to $f_{a}$, the longer is the duration of the parametric resonance for $y_{k}$. Section A.1 discusses this point with further details.
As a result, we can posit an approximate condition to ensure that $U(1)_{\textsc{pq}}$ is not restored due to parametric resonance: for the model of warm inflation considered in this section, this condition is $f_{a}^{\text{(inf)}}/f_{a}\lesssim 700$. As clear in Fig. 4, when this condition is satisfied, the amplification $\delta\theta_{\text{(par.res.)}}$ of the power spectrum of the misalignment angle remains $\mathcal{O}(10^{-3})$, ensuring the validity of our linearised analysis. This condition is model dependent: the details of the saxion-inflaton coupling and of the model for inflation and reheating, affect this quantitative threshold. Depending on the complexity of the dynamics, and on the possibility (for cold inflation) that the oscillations of the inflaton at preheating might act as additional source of parametric resonance, this condition to prevent string formation might be a bit less sharp than in Fig. 4. At any rate, also for the models of cold inflation that we consider in Appendix B we find that the condition $f_{a}^{\text{(inf)}}/f_{a}\lesssim\mathcal{O}(100)$ is an excellent proxy to avoid the formation of axion strings.
@@
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\left(\frac{90M_{\textsc{p}}^{2}}{\pi^{2}g_{\star}}\right)^{1/4}\varepsilon_{i}}$ $<\frac{f_{a}}{H_{I}^{1/2}}$ avoidance of PQ restoration at reheating, Eq. 20 (23a)
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\frac{2.0\cdot 10^{-6}}{\,\text{GeV}^{7/5}}}$ $>\frac{f_{a}}{H_{I}^{12/5}}$ standard isocurvature for axion DM, Eqs. 6 and 7 (23b)
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\frac{\,\text{GeV}^{7/6}}{7.5\cdot 10^{-7}}}\frac{\xi}{\lambda_{\Phi}}$ $>f_{a}^{7/6}$ isocurvature bound for axion DM with $f_{a}^{\text{(inf)}}$, Eq. 12 (23c)
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\frac{4\sqrt{3}}{M_{\textsc{p}}}}\frac{\xi}{\lambda_{\Phi}^{1/2}}$ $<\frac{1}{H_{I}}$ saxion energy density not exceeding $\rho_{\text{tot}}$, Eq. 21 (23d)
$\frac{1}{{\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}M_{\textsc{p}}^{1/2}}}\xi$ $<\frac{1}{H_{I}^{1/2}}$ unitarity bound on $\xi$ from graviton-mediated $\Phi\Phi\to\Phi\Phi$ (23e)
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\frac{4\sqrt{6}}{M_{\textsc{p}}}}\frac{\xi}{\lambda_{\Phi}^{1/2}}$ $\lesssim\frac{1}{H_{I}}$ EFT validity ($2\xi(f_{a}^{\text{(inf)}})^{2}R$ sub-leading wrt $\tfrac{1}{2}M_{\textsc{p}}^{2}R$) (23f)
${\color[rgb]{0.5,0.5,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.65}\mathcal{O}\left(10^{-2}\right)}\sqrt{\frac{\xi}{\lambda_{\Phi}}}$ $\lesssim\frac{f_{a}}{H_{I}}$ avoiding string formation (Fig. 4) (23g)
All of these conditions are set analytically except the last condition Eq. 23g which is an approximate result of our numerical analysis. The numerical coefficient of the last condition would be $\approx\sqrt{24}/700$ in the warm inflation model we consider but is different in other models, though in general is $\sim 10^{-2}$.
@@
Being equipped with Eq. 23g to determine the formation of axion strings, we can understand analytically the results of the scatter plot in Fig. 3. Let us fix the ratio $\xi/\lambda_{\Phi}$, or equivalently $f_{a}^{\text{(inf)}}/H_{I}$. For a fixed value there is an allowed wedge in the plane $(H_{I},f_{a})$, as marked in Fig. 5 for two choices of $f_{a}^{\text{(inf)}}/H_{I}$ (the blue and green regions). The isocurvature bound with $f_{a}^{\text{(inf)}}$, Eq. 23c, gives the horizontal upper limit of these regions. The condition Eq. 23g to prevent string formation translates into a diagonal upper limit on $H_{I}$ as seen in the right hand sides of the blue and green regions in Fig. 5. As we vary $f_{a}^{\text{(inf)}}/H_{I}$, the envelope of these wedges identifies the boundary marked in orange in Fig. 5. If we denote by $R_{f_{a}}$ the upper limit on $f_{a}^{\text{(inf)}}/f_{a}$ identified in Fig. 4 (in that case, $R_{f_{a}}=700$), this right-hand side boundary of the parameter space “rescued” from isocurvature reads (using Eq. 23c for the isocurvature bound)
$f_{a}>8.0\cdot 10^{15}\,\text{GeV}\left(\frac{H_{I}}{10^{11}\,\text{GeV}}\right)^{12/5}\left(\frac{R_{f_{a}}}{100}\right)^{-12/5}\quad\text{(isocurv.\;and strings can be avoided)}$ (24)
This relation is based on the quadratic approximation for the axion potential; for $f_{a}<10^{12}\,\text{GeV}$, the orange line would remain parallel to the isocurvature line shown in grey in Fig. 5. Finally, there is an upper limit on $f_{a}$ that can be derived by the combination of Eqs. 23f and 23c. This bound reads
$f_{a}<0.16\,M_{\textsc{p}}\left(\frac{H_{I}}{10^{11}\,\text{GeV}}\right)^{-6/7}\left(\frac{\lambda_{\Phi}}{10^{-16}}\right)^{-3/7}\quad\text{(isocurv.\;and EFT break.\;can be avoided)}$ (25)
and is marked in purple in Fig. 5, with a choice of $\lambda_{\Phi}=10^{-16}$ for the figures of the paper (a benchmark choice which selects most of our blue points in Fig. 7). This line also roughly coincides with an anaologous combination of the isocurvature constraint Eq. 23c and the requirement that the saxion VEV $f_{a}^{\text{(inf)}}$ during inflation does not reach Planckian values.
@@
We plot in Fig. 7 the points of our scan for the model of Warm Inflation projected on the 2-dimensional subspaces of the variables $(f_{a},\lambda_{\Phi},\xi)$ (while Fig. 3 shows them in the $(H_{I},f_{a})$ plane). Analogously to that figure, the colour code of the points marks in blue the points where $\delta\theta_{\text{(par.res.)}}\lesssim\mathcal{O}(10^{-3})$ and strings are not formed.
@@
  • We neglect the fluctuations $x_{k}$ of the saxion, which display a similar growth as the axion fluctuations and are then irrelevant for us, as long as $\tfrac{1}{\pi}\delta\theta_{\text{(par.res.)}}$ is safely $\ll 1$. One should consider also that the saxion, while passing over the wine bottle potential across the origin, might stop in the angle $\theta_{i}$ (as in the top-left panel of 10) or in $\theta_{i}+\pi$ (as in Fig. 9). This only affects the initial value of the misalignment angle, which we assume to be the right one to yield the axion as DM. We do not need to worry instead that the final angular position ($\theta_{\text{i}}$ or $\theta_{\text{i}}+\pi$) might vary across different spatial patches, depending on the local quantum fluctuations of the saxion. The reason is that, for the “rescued” region of Fig. 6, the saxion crosses the origin less than 10 times, and in that region the relative quantum fluctuations of the saxion $(H_{I}/(2\pi))/f_{a}$ are much smaller than $1/10$, implying that the saxion stops in the same field position across the whole observable universe.
@@
    The decay of the saxion depends on the content of the PQ sector. A model-independent channel is the decay to axions, proportional to $m_{\Phi}^{3}/\langle\phi\rangle^{2}\sim\lambda_{\Phi}^{3/2}f_{a}$ Eroncel:2024rpe. Another channel that is generically present is the decay to SM gluons, also proportional to $m_{\Phi}^{3}/\langle\phi\rangle^{2}$. Other decay channels, as e.g. to KSVZ fermions (when kinematically allowed, see Section A.2), can deplete the saxion energy density into the SM plasma. For the “safe” blue points of the scan shown in Fig. 3, we find that the decays to axions or gluons happen around a similar temperature (due to the same parametric scaling of their decay rates) around $T\sim 10^{2-9}\,\text{GeV}$.
    For about half of the points, the decay occurs after the domination by the saxion of the energy budget of the Universe. In this case, we might worry about isocurvature perturbations on CMB scales. Saxion perturbations on large scales are tiny at early times (the saxion being a massive degree of freedom during inflation Linde:1990flp; Kolb:2023ydq; Racco:2024aac). They inherit the adiabatic perturbations on large scales from the spatial dependence of the Hubble rate: there is effectively a single clock for curvature perturbations. We find then no general issue with the saxion inducing a matter-dominated epoch, given that it decays way before BBN in the parameter space of interest to us.